OpenStudy (medina13):

Please help me :). Will fan. Use Cramer’s rule to solve the system below. Show your work. 2x+4y=6 -3x+y+5

OpenStudy (arthur326):

This is really just about plugging in numbers. The following link lists the general formula for the solutions to a linear 2-by-2 system like yours. https://en.wikipedia.org/wiki/Cramer%27s_rule#Explicit_formulas_for_small_systems Let me know if you're having trouble with any of those expressions.

OpenStudy (medina13):

So i looked at that and it confused me even more! @arthur326

OpenStudy (arthur326):

Well it gives a general 2 by two system, in the same form as the one you have, and it says what \(x\) and \(y\) have to be, which are your solutions.

OpenStudy (arthur326):

Can you point out where you get confused? For example, is it where it mentions determinants, or earlier? Otherwise I don't know what you need help with.

OpenStudy (medina13):

The whole thing. i never been good at this kind of stuff. Do i need like step by step. Or something. because i have no idea how to do this.

OpenStudy (arthur326):

Okay, let's walk through this then.

OpenStudy (medina13):

okay sounds good

OpenStudy (arthur326):

The page says to consider the system \[a_1x+b_1y=c_1\] \[a_2x +b_2 y = c_2.\] What are all the \(a\)'s, \(b\)'s and \(c\)'s in your case?

OpenStudy (medina13):

2 4 6

OpenStudy (medina13):

3 y 5

OpenStudy (arthur326):

Careful; \(b_2\) is not \(y\).

OpenStudy (medina13):

ohok

OpenStudy (arthur326):

What number is really in front of \(y\)?

OpenStudy (medina13):

there isnt one ? or is it 3

OpenStudy (arthur326):

It's definitely not 3. Think about it: the expression \(7y\) for example, means there are 7 \(y\)'s, and the expression \(2y\) mean there are 2 \(y's\). If we just have \(y\), how many \(y's\) do we have?

OpenStudy (medina13):

1 ? or 5 ?

OpenStudy (arthur326):

1 :)

OpenStudy (medina13):

so y turns into 1 ?

OpenStudy (arthur326):

\(y\) is the same thing as \(1y\), so \(b_2=1\). Does that make sense?

OpenStudy (medina13):

so what

OpenStudy (medina13):

some

OpenStudy (arthur326):

Well we know what all the \(a\)s, \(b\)'s and \(c\)'s are, and if you look at the page, it says what \(x\) and \(y\) are in terms of the \(a\)s, \(b\)'s and \(c\)'s. For example, \[x=\dfrac{c_1b_2-b_1c_2}{a_1b_2-b_1a_2}.\] Now you just plug in the numbers.

OpenStudy (medina13):

So how do i plug the numbers ?

OpenStudy (arthur326):

Look at \(c_1b_2\). Since \(c_1 = 6\), and \(b_1 = 1\), the term \(c_1b_2\) is \(6\cdot 1 = 6\).

OpenStudy (arthur326):

(Sorry I meant \(b_2 = 1\).)

OpenStudy (wolf1728):

Here's a helpful page http://www.1728.org/cramer.htm and let's work it out.

OpenStudy (medina13):

I never been so confused !

OpenStudy (wolf1728):

ax + by = e cx + dy = f 2x+4y=6 -3x+y+5 a=2 b=4 e= 6 c=-3 d=1 f=5 denom = |a b| |cd| = a*d -c*b denom = |2 4| |-3 1| = 2*1 --3*4 denom =14

OpenStudy (wolf1728):

x =|e b| |f d| divided by denominator x = |6 4| |5 1| divided by 14 x = (6*1 - 5*4) / 14 x = (6 -20) / 14 x = -14/ 14 x = -1

OpenStudy (wolf1728):

-1 is correct (I already solved the equations)

OpenStudy (medina13):

WOW thank you so much, i understand all of this now !

OpenStudy (medina13):

Wait , so i write down all of that ?

OpenStudy (wolf1728):

I'm still working on "y" but when you use Cramer's rule, you would write all of that down.

OpenStudy (medina13):

ok

OpenStudy (medina13):

and thats all that has to be done ?

OpenStudy (wolf1728):

Yes, but I'm still working on "y" THEN you'll be finished and y does equal 2

OpenStudy (medina13):

okay y = 2

OpenStudy (wolf1728):

y= |a e| |c f| divided by denom y= |2 6| |-3 5| divided by denom y = (2 *5 --3*6) / 14 y = (10 + 18) / 14 y = 28 / 14 y = 2 YAY it's just that simple LOL

OpenStudy (wolf1728):

My mistake was in the definition of "y" (in my first calculation) which I think I'll delete.

OpenStudy (medina13):

your amazing lol

OpenStudy (wolf1728):

Thank you and I think I'll have to change my page for that correction. http://www.1728.org/cramer.htm (That page explains it rather well, except for the definition of "y")

OpenStudy (medina13):

thank you !

OpenStudy (wolf1728):

u r welcome